OpenStudy (anonymous):

recent exam question I need help on use partial fractions to rewrite this integral as the sum of 9 integrals then stop do not evaluate any of the 9 integrals int(2/((x-5)^3(x^2+3)^2(x^2+x+9)))dx

OpenStudy (anonymous):

\[\frac{2}{(x-5)^3(x^2+3)^2(x^2+x+9)}\]

OpenStudy (anonymous):

\[\frac{A}{(x-5)} + \frac{B}{(x-5)^2} + \frac{C}{(x-5)^3} ...\]

OpenStudy (anonymous):

Don't know how to write the last 6 parts, but 4 come from (x^2+3)^2 and 2 come from (x^2+x+9)

OpenStudy (anonymous):

Once you're done setting that up, you set up an equation and solve for A,B,C,D,E,F,G,H, and I. May this problem die in a fire. Lols.

OpenStudy (anonymous):

\[\frac{-44 x-87}{53508 \left(x^2+3\right)^2}-\frac {17 (23 x-333)}{116861472 \left(x^2+3\right)}-\frac{2 (x+6)}{59319 \left(x^2+x+9\right)}+\frac {168913}{4557597408 (x-5)}-\frac{34}{521703 (x-5)^2}+\frac{1}{15288 (x-5)^3} \] No one will expect you to do by hand, unless torture is the objective. Notice that any terms in the denominator that does not have real roots. Need a term of the form U x + V once or more depending of the power of the denominator if the term that does not have real roots.

OpenStudy (anonymous):

\[ \frac{-44 x-87}{53508 \left(x^2+3\right)^2}-\frac {17 (23 x-333)}{116861472 \left(x^2+3\right)}-\frac{2 (x+6)}{59319 \left(x^2+x+9\right)}+\\\frac {168913}{4557597408 (x-5)}-\frac{34}{521703 (x-5)^2}+\frac{1}{15288 (x-5)^3} \]